OpenStudy (anonymous):

Find the least common denominator for these two rational expressions. n^2/n^2+4n+4 and -3/n^2+7n+10

OpenStudy (anonymous):

Helppp anyone!!

OpenStudy (anonymous):

n^2+4n+4=n^2+2n+2n+4=n(n+2)+2(n+2)=(n+2)(n+2) n^2+7n+10=n^2+5n+2n+10=?

OpenStudy (anonymous):

so the answer is (n+2)?

OpenStudy (anonymous):

helppppppp

OpenStudy (anonymous):

correct.